One factory claims that its remote control carts last an average of 320 hours.
To verify this average, a sample of 15 carts is taken. If the calculated value falls within a symmetrical range from the 90% average, the factory is satisfied.
What conclusion would the company obtain according to this sample? The sample
were the following data:

320, 321, 311, 313, 310, 300, 302, 312, 310, 275, 310, 287, 293, 300, 295
Answer #1

The computations are done in the following table here:

X (X - Mean(X))^2
320 258.14
321 291.27
311 49.94
313 82.20
310 36.80
300 15.47
302 3.74
312 65.07
310 36.80
275 837.14
310 36.80
287 286.74
293 119.54
300 15.47
295 79.80

The sample mean here is computed as:

\bar X = \frac{\sum X_i}{n} = \frac{320 + 321 + ... + 295}{15} = 303.93

The sample standard deviation is now computed here as: (from second column)

s = \sqrt{\frac{\sum (X_i - \bar X)^2}{n-1}} = \sqrt{\frac{2214.93}{15-1}} = 12.5781

For 90% confidence level, and n - 1 = 14 degrees of freedom, we have from t distribution tables here:
P( t14 < 1.761) = 0.95

therefore, due to symmetry, we have here:
P( - 1.761 < t14 < 1.761) = 0.9

Therefore the confidence interval here is computed as:
\bar X \pm t^*\frac{s}{\sqrt{n}}

303.93 \pm 1.761*\frac{12.5781}{\sqrt{15}}

303.93 \pm 5.72

(298.21, 309.65)

As the given value of 320 does not lie in the above computed confidence interval, therefore the factory is not satisfied here.